OpenStudy (anonymous):

someone please help with thiss .... 5g+4(-5+3g)=1-g

OpenStudy (mef4.0):

distribute the 4 so that you get 5g-20+12g = 1-g, then combined like terms on the left side, then add 1g to both sides of equation so that all the g's end up on left side, then solve for g

OpenStudy (anonymous):

5g+4(-5+3g)=1-g 5g-20+12g=1-g 17g-20=1-g 17g-21=-g -21=-18g then solve for g